# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

### v1

@@ Abstract.
The emergence of AI-augmented Data Processing Systems (DPSs) has introduced powerful semantic operators that extend traditional data management capabilities with LLM-based processing. However, these systems face fundamental reliability (a.k.a. trust) challenges, as LLMs can generate erroneous outputs, limiting their adoption in critical domains. Existing approaches to LLM constraints—ranging from user-defined functions to constrained decoding—are fragmented, imperative, and lack semantics-aware integration into query execution. To address this gap, we introduce Semantic Integrity Constraints (SICs), a novel declarative abstraction that extends traditional database integrity constraints to govern and optimize semantic operators within DPSs. SICs integrate seamlessly into the relational model, allowing users to specify common classes of constraints (e.g., grounding and soundness) while enabling query-aware enforcement and optimization strategies.
In this paper, we present the core design of SICs, describe their formal integration into query execution, and detail our conception of grounding constraints, a key SIC class that ensures factual consistency of generated outputs. In addition, we explore novel enforcement mechanisms, combining proactive (constrained decoding) and reactive (validation and recovery) techniques to optimize efficiency and reliability. Our work establishes SICs as a foundational framework for trustworthy, high-performance AI-augmented data processing, paving the way for future research in constraint-driven optimizations, adaptive enforcement, and enterprise-scale deployments.
@@ 1. Introduction
Large Language Models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Shankar et al., 2024a; Liu et al., 2025; Anderson et al., 2025; Lu et al., 2025). We broadly refer to these systems as AI-augmented Data Processing Systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. For these systems to be fully adopted, guardrails must be placed around semantic operators.
@@
LLM constraints—also known as assertions or guardrails—have been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related area of work involves constrained decoding, where constraints are applied to the LLM during its generation phase to ensure the output matches a specified structure (Beurer-Kellner et al., 2023; Dong et al., 2024b; guidance, 2022; OpenAI, 2024b; Willard and Louf, 2023; Zheng et al., 2024). While these existing approaches have proven their utility in enhancing the reliability of LLM outputs, they are imperative, fragmented, and lack awareness of operator semantics, limiting their usability and efficiency.
@@
To address these limitations, we introduce Semantic Integrity Constraints (SICs) as a unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to traditional integrity constraints, SICs make the specification of constraints for predominant use cases easy and more advanced cases possible. They are declarative, allowing for a rich set of enforcement and optimization techniques, such as constraint pushdown, constrained decoding, fused evaluation, and adaptive retries. In addition, SICs are expressed in the same language that defines data processing and integrate seamlessly into the relational model. These design decisions are in line with previous efforts that argue for statement and solution specifications on LLM-based system components (Stoica et al., 2024) as well as an (extended) relational model in AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Madden et al., 2024; Lu et al., 2025).
The design of the right interface for SICs and the optimization of their enforcement present many exciting research opportunities and challenges. Different declarative interfaces have been proposed for semantic queries, with the majority adopting the relational model (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025). Although the community has not yet converged on an exact query language, we believe that the language for declaring semantic queries and SICs must be developed in tandem to maintain consistency and tight integration. Moreover, the constraint interface must enable users to declare SICs for the majority of use cases in a simple manner that does not require a deep understanding of the underlying implementation details. Towards this goal, the community will need to conduct user surveys to ensure that design decisions are grounded in real application requirements and do not introduce leaky abstractions. Lastly, optimizing SIC enforcement will require integrating modern techniques in LLM assertions and constrained decoding with traditional database query optimization methods. To meet ever-growing performance demands, novel approaches and extensions to existing solutions must also be developed. These solutions have the potential to accelerate the adoption of AI-augmented solutions in enterprise applications.
This paper presents our vision for SICs and sketches a general SIC framework for relational data processing with unstructured data. Section 2 introduces our proposed constraint interface, followed by Section 3, a categorization of SICs that addresses the most common use cases for LLM constraints. In Section 4, we discuss the enforcement of SICs, detail various optimizations, and highlight multiple promising research directions. Section 5 focuses on grounding constraints, a class of SICs that we are implementing in VectraFlow (Lu et al., 2025), an AI-augmented data flow system developed at Brown University. Finally, we outline several interesting open problems in Section 6.
2. Constraint Interface
We propose an interface that adopts the relational model and enables users to define SICs with the same declarative language used for semantic queries. This approach aligns with traditional methods for expressing integrity constraints, where users declare the logic of a constraint without detailing its physical implementation, allowing the DPS to choose the most efficient way to enforce SICs.
The majority of AI-augmented DPSs provide data flow interfaces, where users chain operators to define a query (Liu et al., 2025; Patel et al., 2024; Anderson et al., 2025). These systems offer a pandas-like interface in Python that exposes semantic operators separately from standard operators. This logical separation prevents declaring both semantic and standard predicates or functions within the same operator. However, recent work has introduced piped data flow syntax to SQL, enhancing its flexibility and extensibility (Shute et al., 2024). This new SQL dialect provides a basis for a unified interface between semantic and traditional queries, while also offering integrity constraints as first class citizens. Even though we design this interface for relational AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025), we expect that it can extend to DPSs with other data models too (Shankar et al., 2024a; Anderson et al., 2025).
@@ 2. Constraint Interface
LABEL:lst:query shows an example query that incorporates semantic functionality into pipe SQL syntax. We first describe the high-level query and then discuss the SIC-specific code later in this and the following sections. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr, biological sex, and age. The query first canonicalizes the sex attribute’s value into “Male” or “Female” (line 2) because the value could have been in another representation, e.g., “M” or “F”. The SET expression is defined using a string prefixed with the letter “p”. This expression declares a prompt string literal (p-string), which specifies functions or predicates that the DPS evaluates with an LLM. Similar to Python’s formatted string literals (f-strings), p-strings can refer to other attributes in the query within curly braces. Lines 5 to 7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Because the EXTEND operator effectively invokes a semantic map to project new attributes, each projected attribute declares a name and type using the AS keyword. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 9 to 10). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the data flow (line 15). Further note that the order of operations specified in the query does not determine the execution order. For example, the DPS should execute the filter on line 9 before summarizing the patient’s medical history on line 8 to avoid excessive computation.
@@
Crucially, our constraint interface extends the ASSERT operator provided by pipe SQL syntax, allowing users to declare SICs anywhere in the data flow. In the example query, assert operators are used in lines 4 and 11 and onward. These operators specify predicates that can refer to all attributes and operators defined upstream. This design enables the assert operator to validate values produced by attribute-generating semantic operators (e.g., maps, aggregates) as well as the presence or absence of tuples from nonattribute-generating operators (e.g., filters, joins, top-).
Since the predicates of the declared assert operators may not be satisfied, the DPS can retry semantic operators in the case of failure. For each semantic operator, the user can specify a retry threshold, which is the maximum number of times the DPS can retry an operator with the same input. The user can also set a retry threshold that applies to all semantic operators in the data flow by default. In LABEL:lst:query, the retry threshold for the SET operator is defined in line 3, and all other operators use the default threshold.
The user can also specify a failure mode (CONTINUE, DROP, or ABORT) for each assert operator, which determines how the DPS handles the failed input tuple(s). Similar to the retry threshold, the user can set a default failure mode. In the example query, line 4 defines the CONTINUE failure mode on the assert operator, while all other assert operators use the default. If a tuple does not satisfy the assert operator’s predicate, then the CONTINUE failure mode continues to propagate it downstream, while DROP drops it. Note that CONTINUE and DROP correspond to the (soft) Suggest and (hard) Assert constructs in DSPy Assertions (Singhvi et al., 2024), respectively. Lastly, ABORT aborts the entire query if a tuple does not satisfy the predicate. While the semantics of aborting are straightforward in a transactional context, in a data flow context, this means that the system will not only drop the tuple that violated the SIC but also discard all of its dependent tuples, such as those that would be in the same aggregation window. In any case, the DPS will log the appropriate errors and warnings when a tuple fails a SIC or when tuples are dropped to notify users of failures.
3. Constraint Classes
@@
We categorize SICs into different classes (Table 1) that are designed to capture the predominant use cases for LLM constraints:
@@
| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHRs |
| Grounding | Value is grounded in the source values | Extracted results or summaries are factual w.r.t. original EHRs |
@@
Domain Constraints. In the context of SICs, domain constraints remain similar to their traditional counterpart, checking whether an attribute’s value is from a specified domain. For example, a domain constraint can enforce that a value is of a certain data type or structure, within a numerical range, from a finite set of possible values, or less than a certain length. (Type definitions of generated attributes from semantic operators implicitly specify domain constraints.) In the example query from LABEL:lst:query, the first conjunct in line 4 and the conjunct in line 14 explicitly define domain constraints, restricting the sex to either “Male” or “Female” and the summary’s length to less than 1000 characters, respectively. This class of SICs comprise the majority of constraint use cases (Liu et al., 2024). Importantly, the DPS can enforce domain constraints programmatically without relying on an LLM validator, providing correctness guarantees on their evaluation.
Inclusion and Exclusion (IE) Constraints. IE constraints determine if particular terms, content, topics, or abstract features like tones and styles are included or excluded from an attribute’s value. For instance, these constraints ensure that an attending physician’s name is present in an EHR, while avoiding the inclusion of PII. Users can specify these constraints with a set of values that indicate what must be included or excluded. This set can be derived from a fixed collection or another attribute’s value. The user can opt for checks based on exact keyword matches or semantic similarity, even using regular expressions for matches on emails or phone numbers. Users can also define IE constraints with a p-string for content, topics, tones, or styles, as shown in line 13 of the example query, which excludes test results from the medical history summary. In this case, the DPS employs an LLM to enforce the constraint, while lightweight models can be used to handle PII or toxic content checks, such as NER models, LSTMs, or feature-based classifiers.
Grounding Constraints. These constraints target both extractive and abstractive use cases, where users want to check whether the output values of semantic operators are grounded in the source values. For instance, line 11 from the example query ensures that the extracted physical exam and lab results appear verbatim in the original EHR (extractive use case). Similarly, line 12 validates that the generated summary of the patient’s medical history is factually consistent with the EHR (abstractive use case). We discuss grounding constraints in greater detail in Section 5.
@@
Soundness Constraints. These constraints verify whether the underlying LLM’s reasoning process is logically sound in that its premises are true and its reasoning steps are valid. This class of constraints targets nonattribute-generating semantic operators and semantic transformations from one or more values to a new value. For such cases, it is nontrivial to simply evaluate predicates on the operator’s output without invoking an LLM validator to redo the task. Instead of this similarly error-prone approach, the DPS uses an LLM validator to check the underlying LLM’s reasoning process, e.g., chain-of-thought (CoT) (Wei et al., 2022). For example, consider the semantic filter (defined on lines 9 and 10), which selects tuples where the patient is predicted to have sepsis based on their medical condition. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to outline sepsis symptoms and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator to ensure that the premises are grounded in the input and that the reasoning steps are valid. This step is defined with the assert operator in line 15. Similarly, a soundness constraint is specified in the second conjunct of line 4, which ensures that sex is canonicalized correctly. If the underlying LLM already performs well in a closed-book setting and the retrieved documents do not provide enough context to produce a valid answer, another approach is to have the LLM generate its CoT first and then verify each proposed premise with the input tuple(s) or documents retrieved afterwards. The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023). This is an avenue for future research.
@@
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given semantic operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query; e.g., the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than their social history. Relevance constraints are evaluated by passing the operator’s prompt, input, and output to an LLM validator for verification.

### v2

@@ Abstract.
AI-augmented data processing systems (DPSs) integrate large language models (LLMs) into query pipelines, allowing powerful semantic operations on structured and unstructured data. However, the reliability (a.k.a. trust) of these systems is fundamentally challenged by the potential for LLMs to produce errors, limiting their adoption in critical domains. To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries. SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both proactive and reactive enforcement strategies.
We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems. Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs. To guide and evaluate the vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
@@ 1. Introduction
Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). We broadly refer to these systems as AI-augmented data processing systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. Without guardrails, it will be difficult to fully adopt these systems.
@@
LLM constraints—also known as assertions or guardrails—have recently been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023; TruLens, 2021; Zheng et al., 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related approach is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
@@
We propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to their traditional counterparts, SICs make defining constraints for predominant use cases easy and more advanced cases possible, aligning with recent work that argues for specifications on LLM-based system components (Stoica et al., 2024).
| Success Criterion | Description |
| Expressiveness | SICs support a range of constraint types—including domain, grounding, and soundness—through a declarative interface (Section 3) |
| Executable Semantics | Each constraint has precise execution semantics—including what is checked, when it is enforced, and what actions are taken on failure—ensuring semantically consistent system behavior and predictable outcomes for users (Section 4) |
| System Integration | SICs integrate with query interfaces, optimizers, and execution engines in AI-augmented DPSs (Section 2) |
| Performance Awareness | Enforcement strategy is selected based on cost and reliability across proactive or reactive methods (Section 4.2) |
| Enterprise-Scale Applicability | SICs support observability, reuse, and auditability across workflows, with mechanisms for conflict detection and policy-driven enforcement at organizational scale (Sections 5 and 6) |
This paper presents our vision for SICs and sketches a general framework for integrating them into AI-augmented DPSs. To evaluate the feasibility and impact of this vision, we define a set of design goals and success criteria: (1) expressiveness, to capture a broad range of common constraint types; (2) executable semantics, with well-defined runtime behavior; (3) system integration, via seamless support within data flow and query processing pipelines; (4) performance-aware enforcement, balancing reliability and cost; and (5) enterprise-scale applicability, including observability, constraint reuse, conflict detection, and auditability. These criteria, outlined in Table 1, were defined through our partnership with the Rhode Island Hospital. They guide our design and are reflected in the structure of the paper. Section 3.1 introduces the SIC interface, while Section 3.2 categorizes SICs across key enforcement classes aligned with real-world LLM use cases. Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms. Section 5 discusses observability components. Finally, Section 6 presents open challenges in large-scale constraint management.
2. System Model
We begin with an overview of the system architecture (Figure 1) for supporting SICs and defer a more detailed discussion of the core components to later sections. First, the user expresses their semantic query along with SICs via the query interface (Section 3.1). This interface forwards the query’s text to the query parser, which produces a corresponding logical query plan. Next, the logical optimizer (Section 4.1) applies logical rewrite rules to the initial plan, yielding an optimized logical plan. This plan is then passed to the physical optimizer (Section 4.2), which considers different physical implementations of the plan and selects the optimal one. The resulting optimized physical plan is then executed by the execution engine. To enable performance monitoring, an observability stack is integrated with the execution engine (Section 5). Finally, the constraint store recommends and analyzes SICs (Section 6).
System architecture for SICs. Orange circles and turquoise diamonds represent semantic operators and SICs, respectively. Green squares indicate annotations that describe the physical implementation of each node.
3. Constraint Specification
We suggest defining SICs declaratively, allowing users to focus on constraint logic while the optimizer determines the best execution method. We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing. User surveys should be conducted to ensure usability.
3.1. Interface
To provide a unified interface for queries and SICs, users declare constraints in the query language. Specifically, we present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs. Though, SICs are compatible with any declarative AI-augmented DPS. The interface allows users to specify two types of information to the DPS for each constraint: (1) what the constraint is and (2) how the DPS should react when the constraint is violated.
Declaring Constraints. SICs specify predicates either on generated tuple attributes from semantic operators or on the semantic operators themselves. We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case. The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT <attribute>, followed by a keyword that refers to the constraint class (Section 3.2) and any necessary parameters. A constraint on an operator is similar: ASSERT <operator>. Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
Retry Thresholds. Users can define a retry threshold for each operator. When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached. To specify a retry threshold, users write RETRY <threshold> after the operator.
Failure Modes. Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with operators and specify what the DPS should do when a failing output reaches the retry threshold. With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream. IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window). Lastly, ABORT aborts the entire query. To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
@@ 3.1. Interface
End-To-End Example. LABEL:lst:query shows an example semantic query with various SICs. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr and dob. The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format. The expression is defined using a string prefixed with the letter “p”, which declares a prompt string. Prompt strings specify natural language functions or predicates over attributes and are evaluated using LLMs. Line 4 computes the patient’s age using standard SQL functions. Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 10-11). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 16).
The first constraint is declared on line 3, and the remaining constraints are specified on line 12 and onward. If the summarization step for med_hist_sum (line 8) fails its constraints (lines 13-15), the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraints (line 9). Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used. We explain each constraint in more detail in the next section.
@@
3.2. Constraint Classes
@@
We categorize SICs into different classes (Table 2) that are designed to capture the predominant use cases for LLM constraints. We motivate each constraint class using concrete examples from the EHR query (among others) and briefly outline their implementation.
Domain Constraints. In many queries, users want the generated attributes to belong to a specified domain. This includes use cases such as enforcing that a value is of a certain data type or structure, from a set of values, or less than a certain length. For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values. We extend traditional domain constraints to semantic data processing. In LABEL:lst:query, lines 3 and 15 specify domain constraints, restricting the dob to the regex representation of YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively. The type definitions (e.g., STRING) of generated attributes implicitly define domain constraints. This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
Inclusion and Exclusion (IE) Constraints. Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles. For example, given a semantic map that summarizes an EHR, users may want to exclude all patient PII from the summary while including the names of attending physicians. IE constraints address this class of predicates. In the example query, line 14 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary. When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match. Users can also specify IE constraints using strings, regular expressions, or even a collection of them (e.g., a set of toxic words to exclude from the output). In such cases, deterministic methods are used to evaluate matches.
Grounding Constraints. Semantic operators, such as maps or aggregates, can extract data or summarize from their inputs. However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data. Grounding constraints enable users to enforce such invariants on generated attributes. For example, line 12 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR. Line 13 asserts that the summarized med_hist_sum is factually consistent with the source EHR. In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Formally, if a grounding constraint is defined on an attribute, the DPS will recursively enforce grounding constraints on all ancestor attributes in its lineage. Thus, each constraint only needs to compare its operator’s input and output to ensure grounding in the source data. A grounding constraint is implemented in one of two ways, depending on whether the specified attribute is generated via an extractive (map) or abstractive (map or aggregate) operator. In the extractive case, the constraint is implemented deterministically and ensures that the operator’s output is precisely contained in its input. In the abstractive case, the constraint is implemented stochastically (e.g., with an LLM or a fact-checking model (Tang et al., 2024)) and ensures that the operator’s summarized output is factually consistent with the input. Users can annotate relevant operators with the EXTRACTIVE (lines 5-7) or ABSTRACTIVE (line 8) keywords to inform the DPS of their semantics; otherwise, the DPS infers these annotations by reasoning about the operator’s prompt string with an LLM.
@@
| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHR summaries |
| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. original EHRs |
@@
Soundness Constraints. To check whether the reasoning of operators themselves is logically sound, users can declare soundness constraints. Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-. To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning process by modifying the operator to output its reasoning steps, e.g., chain-of-thought (CoT) (Wei et al., 2022). Line 16 declares a soundness constraint on the sepsis filter. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to describe the symptoms of sepsis and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid. The DPS can also use more sophisticated CoT prompting and verification techniques, which use structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
@@
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query. For example, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history. These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.